# Robust topological invariants of timelike circular orbits for spinning test particles in black hole spacetimes

### v2

@@ I Introduction
Circular orbits in black-hole spacetimes are essential tools for probing strong-field gravity, underpinning key astrophysical phenomena such as accretion disks, black-hole shadows, and gravitational-wave emission from extreme mass-ratio inspirals (EMRIs). Their study provides deep insight into the structure of spacetime and the behavior of matter in extreme environments.
@@
Although significant progress has been made in understanding null circular orbits, the case of timelike circular orbits (TCOs) is more intricate, owing to their dependence on orbital parameters such as angular momentum. Earlier studies established that, for neutral particles with fixed angular momentum, TCOs generically arise in stable–unstable pairs Wei:2022mzv ; Ye:2023gmk ; this pairing is topologically enforced and independent of black-hole parameters. Building on this, several further studies have been carried out Shahzad:2024oxb , including the case of charged particles. It has been shown that the electric charge of the particle indeed alters the topological properties of its circular orbits Ye:2024sus ; Song:2025vhw .
@@
This gap in understanding motivates our central research question: does the particle’s spin, like its charge, alter the topological properties of circular orbits? Although the behavior of effective potentials for spinning particles and the existence of circular orbits have been studied since the pioneering work of Tod, de Felice, and Calvani Tod:1976ud , nd spin is known to produce quantitative shifts in orbital parameters Toshmatov:2019bda ; Toshmatov:2020wky , it remains unknown whether it changes the fundamental topological invariants that characterize the global orbital structure. If these invariants are robust against spin, it would imply a deeper level of universality in orbital dynamics, dictated purely by spacetime geometry.
@@
II TOPOLOGICAL APPROACH FOR CIRCULAR ORBITS
In this section, we present the general formalism for a spinning test particle within the Mathisson–Papapetrou–Dixon (MPD) approximation, up to the pole-dipole order Mathisson:1937zz ; Papapetrou:1951pa ; Dixon:1964 ; Dixon:1965:1 ; Dixon:1970:1 ; Dixon:1970:2 ; Dixon:1973:1 ; Dixon:1974:1 ; Dixon:1979 and review the topological approach for studying the properties of circular orbits. We consider a spinning particle moving in a static, spherically symmetric spacetime, with the line element:
@@ II TOPOLOGICAL APPROACH FOR CIRCULAR ORBITS
$\dot{P}^{a}=-\frac{1}{2}R^{a}{}_{bcd}u^{b}S^{cd}\;,$ (2)
$\dot{S}^{ab}=2P^{[a}u^{b]}\;.$ (3)
where $D/d\lambda$ is the covariant derivative along the particle’s trajectory with the affine parameter $\lambda$ given by $D/d\lambda\equiv u^{a}\nabla_{a}$, and $R^{a}_{bcd}$ is the Riemann tensor. The dynamical 4-momentum and kinematical 4-velocity of the particle are denoted by $P^{a}$ and $u^{a}$, respectively, and the anti-symmetric spin tensor is denoted by $S_{ab}$ (with $S_{ab}=-S_{ba}$).
@@
where $s=S/\mathcal{M}$ is specific spin parameter. It should be noted that $s$ can have both negative and positive values depending on the direction of spin with respect to the direction of the orbital angular momentum $L$. While $L$ is conserved, $P_{\phi}$ is not necessarily constant, see Eq. (11). The sign of $s$ relative to $L$ defines co/counter-rotation. In the spherically symmetric case, $L$ can always be chosen to be greater than zero, i.e., $L>0$. Then, when $s<0$, the spin and angular momentum are counter-rotating; when $s>0$, the spin and angular momentum are co-rotating. Toshmatov:2020wky . From the conservation of energy (10) and angular momentum (11), we have
@@
To ensure the mathematical consistency of the derived expressions (e.g., avoiding divergences in $p_{\phi}$, see Eq.(16)) and to maintain the timelike character of the particle’s four-velocity $u^{a}$, see Eq.(8), (thus preserving causality outside the event horizon), we restrict our analysis to cases where the spin magnitude is small enough such that:
@@
where the factor $\frac{1}{\sin\theta}$ is an auxiliary term. This construction is motivated by the need to define a vector field $\phi^{a}$ (see below) that is regular on the entire $(r,\theta)$ plane, including the poles $\theta=0,\pi$, and whose zeros correspond to circular orbits on the equatorial plane. The factor $1/\sin\theta$ ensures the asymptotic behavior of $\phi^{a}$ at the boundaries $\theta=0,\pi$ is universal and independent of $L$ and $s$, which is crucial for a consistent topological classification.
@@
where $w_{i}$ is the winding number of the $i$-th zero point within region $\Sigma$. Alternatively, $W$ can be computed as:
$W=\frac{1}{2\pi}\oint_{I}d\Omega\;.$ (35)
where $I=\partial\Sigma$ is a counterclockwise closed contour enclosing $\Sigma$. We choose the contour $C=\sum_{i}\cup l_{i}$ as shown in Fig. 1.
@@ III Between two neighboring horizons
Figure 3: Representation of the contour $C=\sum_{i}l_{i}$ (which encloses $\Sigma$) on the $(r,\theta)$ plane. The curve $C$ has a positive orientation, marked with the red arrows. The black and blue arrows indicate the approximate directions of the vector $\phi$ at the boundaries.
@@
Ignoring the specific value of $\phi^{\theta}$, the vector $\phi$ points to the right on $l_{3}$.
@@
so $\phi$ points to the left on $l_{1}$.
These results are shown by blue arrows in Fig. 3. Combining the behavior at $\phi$ at $\theta=0$ and $\theta=\pi$, we obtain
@@ IV.1 Asymptotically flat black holes
$\phi_{l_{1}}^{r}(r\to\infty)\sim\frac{M\mathcal{M}}{r^{2}}\;,$ (42)
@@
    Figure 7: The graph of the potential function (46) for $L=1$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
@@
    Figure 8: The graph of the potential function (46) for $L=2$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
@@ IV.2 Asymptotically AdS black holes
    Figure 12: The graph of the potential function (52) for $L\approx 2.526$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively. The red and the blue line represent $s=0$ and $s\approx 0.5$, respectively.
@@
    Figure 13: The graph of the potential function (52) for $L=1.5$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
@@
    Figure 14: The graph of the potential function (52) for $L=3$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
@@ IV.3 Asymptotically dS black holes
indicating at least one unstable TCO for fixed $L$ and $s$. The cosmological horizon in dS spacetime introduces a natural boundary that modifies the topological classification. The guaranteed unstable TCO between the black hole and cosmological horizons may have implications for the global structure of orbits in our expanding universe, particularly for supermassive black holes where this region is astrophysically accessible.
@@
Figure 17: The graph of the potential function (56) for $L=0.3$. The red points and the blue boxes represent the maximum values and the minimum values of (56) for a given value of $s$, respectively.
@@ V Conclusions and discussions
n this work, we demonstrate the power of topological methods in classifying circular orbits, even for spinning particles. The invariance of $W$ under spin perturbations suggests a deep robustness in the orbital structure dictated by the background metric. This finding has important implications for gravitational wave astronomy, particularly for modeling extreme mass-ratio inspirals where the secondary object may possess significant spin. Our main conclusions are:
@@
The spin-independence of the topological winding number implies that the fundamental orbital structure—such as the guaranteed existence of unstable TCOs between horizons or the pairing of stable and unstable TCOs outside—is a property of the spacetime itself, not of the particle’s internal composition. This universality underpins the robustness of gravitational wave signatures from spinning secondaries in EMRIs, suggesting that certain waveform features are more generic than previously assumed.

### v3

@@ I Introduction
Circular orbits in black hole spacetimes are essential tools for probing strong-field gravity, underpinning key astrophysical phenomena such as accretion disks, black-hole shadows, and gravitational-wave emission from extreme mass-ratio inspirals (EMRIs). Their study provides deep insight into the structure of spacetime and the behavior of matter in extreme environments.
@@
Although significant progress has been made in understanding null circular orbits, the case of timelike circular orbits (TCOs) is more intricate, owing to their dependence on orbital parameters such as angular momentum. Earlier studies established that, for neutral particles with fixed angular momentum, TCOs generically arise in stable–unstable pairs Wei:2022mzv ; Ye:2023gmk ; this pairing is topologically enforced and independent of black-hole parameters. Building on this, several further studies have been carried out Shahzad:2024oxb ; Ye:2024sus ; Song:2025vhw , including the case of charged particles. It has been shown that the electric charge of the particle indeed alters the topological properties of its circular orbits Ye:2024sus ; Song:2025vhw .
@@
This gap in understanding motivates our central research question: does the particle’s spin, like its charge, alter the topological properties of circular orbits? Although the behavior of effective potentials for spinning particles and the existence of circular orbits have been studied since the pioneering work of Tod, de Felice, and Calvani Tod:1976ud , and spin is known to produce quantitative shifts in orbital parameters Toshmatov:2019bda ; Toshmatov:2020wky , it remains unknown whether it changes the fundamental topological invariants that characterize the global orbital structure. If these invariants are robust against spin, it would imply a deeper level of universality in orbital dynamics, dictated purely by spacetime geometry.
@@
II Mathisson-Papapetrou-Dixon equations and topological approach for circular orbits
In this section, we present the general formalism for a spinning test particle within the Mathisson-Papapetrou-Dixon (MPD) approximation, up to the pole-dipole order Mathisson:1937zz ; Papapetrou:1951pa ; Dixon:1964 ; Dixon:1965:1 ; Dixon:1970:1 ; Dixon:1970:2 ; Dixon:1973:1 ; Dixon:1974:1 ; Dixon:1979 and review the topological approach for studying the properties of circular orbits. We consider a spinning particle moving in a static, spherically symmetric spacetime, with the line element:
@@ II Mathisson-Papapetrou-Dixon equations and topological approach for circular orbits
$\frac{DP^{a}}{d\lambda}=-\frac{1}{2}R^{a}{}_{bcd}u^{b}S^{cd}\;,$ (2)
$\frac{DS^{ab}}{d\lambda}=2P^{[a}u^{b]}\;.$ (3)
where $D/d\lambda$ is the covariant derivative along the particle’s trajectory with the affine parameter $\lambda$ given by $D/d\lambda\equiv u^{a}\nabla_{a}$, and $R^{a}{}_{bcd}$ is the Riemann tensor. The dynamical 4-momentum and kinematical 4-velocity of the particle are denoted by $P^{a}$ and $u^{a}$, respectively, and the anti-symmetric spin tensor is denoted by $S^{ab}$ (with $S^{ab}=-S^{ba}$).
@@
where $s=S/\mathcal{M}$ is specific spin parameter. It should be noted that $s$ can have both negative and positive values depending on the direction of spin with respect to the direction of the orbital angular momentum $L$. While $L$ is conserved, $P_{\phi}$ is not necessarily constant, see Eq. (11). The sign of $s$ relative to $L$ defines co/counter-rotation. In the spherically symmetric case, $L$ can always be chosen to be greater than zero, i.e., $L>0$. Then, when $s<0$, the spin and angular momentum are counter-rotating; when $s>0$, the spin and angular momentum are co-rotating Toshmatov:2020wky . From the conservation of energy (10) and angular momentum (11), we have
@@
To ensure the mathematical consistency of the derived expressions (e.g., avoiding divergences in $p_{\phi}$, see Eq. (16)) and to maintain the timelike character of the particle’s four-velocity $u^{a}$, see Eq. (8), (thus preserving causal structure outside the event horizon), we restrict our analysis to cases where the spin magnitude is small enough such that:
@@
where the factor $\frac{1}{\sin\theta}$ is an auxiliary term [^4] The choice of the factor $\frac{1}{\sin\theta}$ is not unique; any function that maintains a universal asymptotic behavior at $\theta=0,\pi$ and preserves the correspondence between zeros of $\phi^{a}$ and equatorial circular orbits would yield the same topological conclusions. For instance, a factor like $1/\sin^{2}\theta$ or other powers of $\sin\theta$ could also be used, provided the resulting vector field remains well-behaved at the boundaries.. This construction is motivated by the need to define a vector field $\phi^{a}$ (see below) that is regular on the entire $(r,\theta)$ plane, including the poles $\theta=0,\pi$, and whose zeros correspond to circular orbits on the equatorial plane. The factor $1/\sin\theta$ ensures the asymptotic behavior of $\phi^{a}$ at the boundaries $\theta=0,\pi$ is universal and independent of $L$ and $s$, which is crucial for a consistent topological classification.
@@
where $w_{i}$ is the winding number of the $i$-th zero point within region $\Sigma$.
The topological number $W$ defined in Eq. (34) can be equivalently expressed as a contour integral of the angular function $\Omega$ introduced in Eq. (29). To see this, note that the topological current (30) can be rewritten in terms of a gauge potential $A_{\mu}=\epsilon_{ab}n^{a}\partial_{\mu}n^{b}$. Using the unit vector $n^{a}=(\cos\Omega,\sin\Omega)$, one finds $A_{\mu}=\partial_{\mu}\Omega$. The topological current density then becomes
$j^{0}=\frac{1}{2\pi}(\partial_{1}A_{2}-\partial_{2}A_{1})=\frac{1}{2\pi}(\partial_{1}\partial_{2}\Omega-\partial_{2}\partial_{1}\Omega)\;,$ (35)
which, when integrated over the region $\Sigma$, yields via Stokes’ theorem
$W=\int_{\Sigma}j^{0}d^{2}x=\frac{1}{2\pi}\oint_{\partial\Sigma}A_{\mu}dx^{\mu}=\frac{1}{2\pi}\oint_{\partial\Sigma}d\Omega\;.$ (36)
Here $\partial\Sigma=C$ is a counterclockwise closed contour enclosing $\Sigma$. This derivation shows that Eq. (36) computes the same topological invariant as Eq. (34); it provides a practical means of evaluating $W$ by tracking the winding of the vector field $\phi^{a}$ along the boundary. Below, we choose the contour $C=\sum_{i}\cup l_{i}$ as shown in Fig. 1.
@@ III Between two neighboring horizons
Ignoring the specific value of $\phi^{\theta}$, the vector $\phi$ points to the right on $l_{3}$ (blue arrow).
@@
so $\phi$ points to the left on $l_{1}$ (yellow arrow).
Combining the behavior at $\phi$ at $\theta=0$ and $\theta=\pi$, we obtain
@@ IV.1 Asymptotically flat black holes
$\phi_{l_{1}}^{r}(r\to\infty)\sim\frac{M\mathcal{M}}{r^{2}}>0\;,$ (43)
@@
    Figure 7: The graph of the potential function (47) for $L=1$. The red dots and blue squares represent the maximum values and the minimum values of (47) for a given value of $s$, respectively.
@@
    Figure 8: The graph of the potential function (47) for $L=2$. The red dots and blue squares represent the maximum values and the minimum values of (47) for a given value of $s$, respectively.
@@ IV.2 Asymptotically AdS black holes
    Figure 12: The graph of the potential function (53) for $L\approx 2.526$. The red dots and blue squares represent the maximum values and the minimum values of (53) for a given value of $s$, respectively. The red and the blue line represent $s=0$ and $s\approx 0.5$, respectively.
@@
    Figure 13: The graph of the potential function (53) for $L=1.5$. The red dots and blue squares represent the maximum values and the minimum values of (53) for a given value of $s$, respectively.
@@
    Figure 14: The graph of the potential function (53) for $L=3$. The red dots and blue squares represent the maximum values and the minimum values of (53) for a given value of $s$, respectively.
@@ IV.3 Asymptotically dS black holes
indicating at least one unstable TCO for fixed $L$ and $s$. The cosmological horizon in dS spacetime introduces a natural boundary that modifies the topological classification. The guaranteed unstable TCO between the black hole and cosmological horizons may have implications for the global structure of orbits in an expanding universe, particularly for supermassive black holes where this region is astrophysically accessible.
@@
Figure 17: The graph of the potential function (57) for $L=0.3$. The red dots and blue squares represent the maximum values and the minimum values of (57) for a given value of $s$, respectively.
@@ V Conclusions and discussions
In this work, we demonstrate the power of topological methods in classifying circular orbits, even for spinning particles. The invariance of $W$ under spin perturbations suggests a deep robustness in the orbital structure dictated by the background metric. This finding has important implications for gravitational wave astronomy, particularly for modeling extreme mass-ratio inspirals where the secondary object may possess significant spin. Our main conclusions are:
@@
The spin-independence of the topological winding number implies that the fundamental orbital structure–such as the guaranteed existence of unstable TCOs between horizons or the pairing of stable and unstable TCOs outside the outermost horizon–is a property of the spacetime itself, not of the particle’s internal composition. This universality underpins the robustness of gravitational wave signatures from spinning secondaries in EMRIs, suggesting that certain waveform features are more generic than previously assumed.
@@
Acknowledgement
This work is supported by the Research Start-up Funding of Chengdu University of Technology (Grant No. 10912-KYQD2022-09307).